# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v2

@@ 3.5. Annihilators of quotients $I_{1}/I_{2}$
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}\,$. The annihilator of the quotient $I_{1}/I_{2}$ with $I_{2}\subsetneq I_{1}$ is $I_{2}:I_{1}$; in the remaining case (where $I_{1}/I_{2}=0$) we leave the annihilator undefined. In general, the equation $I_{1}J=I_{2}$ is not solvable and $J=I_{2}:I_{1}$ will not be a solution, but it will be the largest $\mathcal{O}_{v}$-ideal such that $I_{1}J\subseteq I_{2}\,$. However, in the situation appearing in the papers [3, 4], we have $I_{1}=U$ and $I_{2}=UV$ for $\mathcal{O}_{v}$-ideals $U,V$. Then $J=V$ is a solution of $I_{1}J=I_{2}$ and our task is just to compute the largest solution $I_{2}:I_{1}$, which we accomplish in the next theorem.
@@
Take $\mathcal{O}_{v}$-ideals $I$ and $J$ such that $IJ\subsetneq I$. Then the annihilator of $I/IJ$ is
@@
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subsetneq I_{1}\,$. If the annihilator of $I_{1}/I_{2}$ is $\mathcal{M}_{v}\,$, then $I_{1}$ is a principal $\mathcal{O}_{v}$-ideal.
@@
If $I_{1}$ is nonprincipal, then by part 3) of Lemma 2.6 we obtain $vI_{1}=vI_{1}+0^{+}=vI_{1}+v\mathcal{M}_{v}\,$, whence $I_{1}=I_{1}\mathcal{M}_{v}\,$. As $\mathcal{M}_{v}$ annihilates $I_{1}/I_{2}$, it follows that $I_{1}=I_{1}\mathcal{M}_{v}\subseteq I_{2}\,$, which contradicts our assumption. ∎
@@
1) Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subsetneq I_{1}\,$. Then the annihilator of $I_{1}/I_{2}$ is equal to $\mathcal{M}_{v}$ if and only if $I_{1}$ is principal and $I_{2}=I_{1}\mathcal{M}_{v}\,$.
2) Take $\mathcal{O}_{v}$-ideals $I$ and $J$ with $IJ\subsetneq I$. Then the annihilator of $I/IJ$ is equal to $\mathcal{M}_{v}$ if and only if $I$ is principal and $J=\mathcal{M}_{v}\,$.
@@
In order to prove the converse implication, assume that $\mathcal{M}_{v}$ is the annihilator of $I_{1}/I_{2}$. From Lemma 3.16 we infer that $I_{1}$ must be principal, say $I_{1}=(a)$ for some $a\in K$. As $\mathcal{M}_{v}$ annihilates $I_{1}/I_{2}\,$, we have $a\mathcal{M}_{v}=I_{1}\mathcal{M}_{v}\subseteq I_{2}\,$. We cannot have $a\mathcal{M}_{v}\subsetneq I_{2}$ since then $I_{1}=(a)\subseteq I_{2}\,$, which contradicts our assumption. Thus we have $I_{2}=a\mathcal{M}_{v}=I_{1}\mathcal{M}_{v}\,$.
@@
Take $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subsetneq I_{1}\,$.
@@ 3.6. Application to the computation of annihilators in special cases
Throughout, we assume that $2\leq n\in\mathbb{N}$. For a nonzero $\mathcal{O}_{v}$-module $M$ we denote by $\mbox{\rm ann}\,\,M$ the annihilator of $M$.
@@
Take an $\mathcal{O}_{v}$-ideal $I$ and set $J=bI^{n-1}$, where $b\in\mathcal{O}_{v}\,$. Assume that $I\neq IJ$.
@@
1): From Theorem 3.15 we infer that $\mbox{\rm ann}\,I/IJ=J$ if $I$ is a principal $\mathcal{O}(J)$-ideal, and $\mbox{\rm ann}\,I/IJ=J^{\diamondsuit}$ if $I$ is a nonprincipal $\mathcal{O}(J)$-ideal. Assume first that there is $a\in K$ such that $vJ/{\mathcal{G}}(vJ)$ has infimum $va/{\mathcal{G}}(vJ)$ in $vK/{\mathcal{G}}(vJ)$ but does not contain this infimum. Then $vJ/{\mathcal{G}}(vJ)$ is nonprincipal, and the same must hold for $vI/{\mathcal{G}}(vJ)$, so that $I$ is a nonprincipal $\mathcal{O}(J)$-ideal. It follows that $\mbox{\rm ann}\,I/IJ=J^{\diamondsuit}=a\mathcal{O}(J)$. Now assume the remaining case, that is, $vJ/{\mathcal{G}}(vJ)$ is closed in $G/{\mathcal{G}}(vJ)$ and therefore, $J^{\diamondsuit}=J$. In this case we always obtain that $\mbox{\rm ann}\,I/IJ=J$.
@@
2): By part 2) of Proposition 3.17, the annihilator of $I/IJ$ is equal to $\mathcal{M}_{v}$ if and only if $I$ is principal and $J=\mathcal{M}_{v}\,$. Assume first that the latter is the case. Since $J=bI^{n-1}$, it follows that $J$ and $\mathcal{M}_{v}$ are principal. Since $I$ is principal, $vI$ is principal and ${\mathcal{G}}(vJ)={\mathcal{G}}(vI)=\{0\}$, so $vJ/{\mathcal{G}}(vI)=vJ$ is principal and from part 1) we obtain that $\mathcal{M}_{v}=\mbox{\rm ann}\,I/IJ=J=bI^{n-1}$. Since $\mathcal{M}_{v}=J\subseteq I$, we have either $I=\mathcal{M}_{v}=J$ or $I=\mathcal{O}_{v}$ and $J=\mathcal{M}_{v}=(b)$. Since $\mathcal{M}_{v}$ is principal, the former holds if and only if $n=2$ and $b\in\mathcal{O}_{v}^{\times}$, and the latter holds if and only if $n=1$ and $b\notin\mathcal{O}_{v}^{\times}$.
@@
From now on we assume that $a\in\mathcal{O}_{v}\,$. As before, we take $\mathcal{O}\in\mathcal{R}$ with maximal ideal $\mathcal{M}$ and denote the valuation associated with $\mathcal{O}$ by $w$. The following facts are used in [4].
@@
We have that
@@
1) The following holds if $a\mathcal{M}\neq(a\mathcal{M})^{n}$:
$\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\>=\>\left\{\begin{array}[]{ll}(a\mathcal{M})^{n-1}&\mbox{if $\mathcal{M}$ is a principal $\mathcal{O}$-ideal, }\\ (a\mathcal{O})^{n-1}=a^{n-1}\mathcal{O}&\mbox{if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal.}\end{array}\right.$
@@
It remains to show that $J\mathcal{O}(IJ)=J$ is equal to $(a\mathcal{O})^{n-1}$ if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. In this case, 00 is the infimum of $w\mathcal{M}$, not contained in $w\mathcal{M}$. Therefore, $wa^{n-1}\notin wa^{n-1}+w\mathcal{M}=wa^{n-1}\mathcal{M}$ is the infimum of $wa^{n-1}\mathcal{M}$. Hence by definition, $J^{\diamondsuit}=\{c\mid wc\geq wa^{n-1}\}=a^{n-1}\mathcal{O}=(a\mathcal{O})^{n-1}$. This completes the proof of part 1).
2): By assumption, $a\mathcal{M}\neq(a\mathcal{M})^{n}$, which by part 3) of Lemma 3.20 means that $a\in\mathcal{M}$ or $\mathcal{M}$ is a principal $\mathcal{O}$-ideal. If the latter holds, then by part 1), $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=(a\mathcal{M})^{n-1}$. Since $a\mathcal{M}\subseteq\mathcal{M}_{v}\,$, this is equal to $\mathcal{M}_{v}$ if and only if $n-1=1$, $\mathcal{M}=\mathcal{M}_{v}\,$, and $va=0$, where we have again used part 3) of Lemma 3.20. On the other hand, if the latter does not hold, then $a\in\mathcal{M}$ and by part 1), $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=(a\mathcal{O})^{n-1}\subseteq a\mathcal{O}$. The latter is contained in $\mathcal{M}$ but is not equal to $\mathcal{M}$ since $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. Thus $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\subsetneq\mathcal{M}\subseteq\mathcal{M}_{v}$ and consequently, $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\neq\mathcal{M}_{v}\,$. ∎

### v3

@@ 3.5. Annihilators of quotients $I_{1}/I_{2}$
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}\,$. The annihilator of the quotient $I_{1}/I_{2}$ with $I_{2}\subseteq I_{1}$ is the $\mathcal{O}_{v}$-ideal $I_{2}:I_{1}$. In general, the equation $I_{1}J=I_{2}$ is not solvable and $J=I_{2}:I_{1}$ will not be a solution, but it will be the largest $\mathcal{O}_{v}$-ideal such that $I_{1}J\subseteq I_{2}\,$. However, in the situation appearing in the papers [3, 4], we have $I_{1}=U$ and $I_{2}=UV$ for $\mathcal{O}_{v}$-ideals $U,V$. Then $J=V$ is a solution of $I_{1}J=I_{2}$ and our task is just to compute the largest solution $I_{2}:I_{1}$, which we accomplish in the next theorem.
@@
Take $\mathcal{O}_{v}$-ideals $I$ and $J$. Then
@@
Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subseteq I_{1}\,$. If the annihilator of $I_{1}/I_{2}$ is $\mathcal{M}_{v}\,$, then $I_{1}$ is a principal $\mathcal{O}_{v}$-ideal.
@@
If $I_{1}$ is nonprincipal, then $vI_{1}$ is nonprincipal and by part 3) of Lemma 2.6 we obtain $vI_{1}=vI_{1}+0^{+}=vI_{1}+v\mathcal{M}_{v}\,$, whence $I_{1}=I_{1}\mathcal{M}_{v}\,$. As $\mathcal{M}_{v}$ annihilates $I_{1}/I_{2}$, it follows that $I_{1}=I_{1}\mathcal{M}_{v}\subseteq I_{2}\,$, which contradicts our assumption. ∎
@@
1) Take $\mathcal{O}_{v}$-ideals $I_{1}$ and $I_{2}$ with $I_{2}\subseteq I_{1}\,$. Then the annihilator of $I_{1}/I_{2}$ is equal to $\mathcal{M}_{v}$ if and only if $I_{1}$ is principal and $I_{2}=I_{1}\mathcal{M}_{v}\,$.
2) Take $\mathcal{O}_{v}$-ideals $I$ and $J$ with $J\subseteq\mathcal{O}_{v}\,$. Then the annihilator of $I/IJ$ is equal to $\mathcal{M}_{v}$ if and only if $I$ is principal and $J=\mathcal{M}_{v}\,$.
@@
In order to prove the converse, assume that $\mathcal{M}_{v}$ is the annihilator of $I_{1}/I_{2}\,$. From Lemma 3.16 we infer that $I_{1}$ must be principal, say $I_{1}=(a)$ for some $a\in K$. As $\mathcal{M}_{v}$ annihilates $I_{1}/I_{2}\,$, we have $a\mathcal{M}_{v}=I_{1}\mathcal{M}_{v}\subseteq I_{2}\,$. We cannot have $a\mathcal{M}_{v}\subsetneq I_{2}$ since then $I_{1}=(a)\subseteq I_{2}\,$, which contradicts our assumption. Thus we have $I_{2}=a\mathcal{M}_{v}=I_{1}\mathcal{M}_{v}\,$.
@@
Take $\mathcal{O}_{v}$-ideals $I_{1},I_{2}$ with $I_{2}\subseteq I_{1}\,$.
@@ 3.6. Application to the computation of annihilators in special cases
Throughout, we assume that $2\leq n\in\mathbb{N}$. For an $\mathcal{O}_{v}$-module $M$ we denote by $\mbox{\rm ann}\,\,M$ the annihilator of $M$.
@@
Take an $\mathcal{O}_{v}$-ideal $I\subseteq\mathcal{O}_{v}$ and set $J=bI^{n-1}$, where $b\in\mathcal{O}_{v}\,$.
@@
1): From Theorem 3.15 we infer that $\mbox{\rm ann}\,I/IJ=J$ if $I$ is a principal $\mathcal{O}(J)$-ideal, and $\mbox{\rm ann}\,I/IJ=J^{\diamondsuit}$ if $I$ is a nonprincipal $\mathcal{O}(J)$-ideal. Assume first that there is $a\in K$ such that $vJ/{\mathcal{G}}(vJ)$ has infimum $va/{\mathcal{G}}(vJ)$ in $vK/{\mathcal{G}}(vJ)$ but does not contain this infimum. Then $vJ/{\mathcal{G}}(vJ)$ is nonprincipal, and the same must hold for $vI/{\mathcal{G}}(vJ)$, so that $I$ is a nonprincipal $\mathcal{O}(J)$-ideal. It follows that $\mbox{\rm ann}\,I/IJ=J^{\diamondsuit}=a\mathcal{O}(J)$. Now assume the remaining case, that is, $vJ/{\mathcal{G}}(vJ)$ is closed in $G/{\mathcal{G}}(vJ)$ and therefore, $J^{\diamondsuit}=J$. In this case we obtain that $\mbox{\rm ann}\,I/IJ=J$.
@@
2): By part 2) of Proposition 3.17, the annihilator of $I/IJ$ is equal to $\mathcal{M}_{v}$ if and only if $I$ is principal and $J=\mathcal{M}_{v}\,$. Assume first that $I$ is principal and $J=\mathcal{M}_{v}\,$. Since $J=bI^{n-1}$, it follows that $J$ and $\mathcal{M}_{v}$ are principal. Since $I$ is principal, $vI$ is principal and ${\mathcal{G}}(vJ)={\mathcal{G}}(vI)=\{0\}$, so $vJ/{\mathcal{G}}(vI)=vJ$ is principal and from part 1) we obtain that $\mathcal{M}_{v}=\mbox{\rm ann}\,I/IJ=J=bI^{n-1}\subseteq I$. Thus we either have that $I=\mathcal{M}_{v}=J$, or that $I=\mathcal{O}_{v}$ and $J=\mathcal{M}_{v}=(b)$. Since $\mathcal{M}_{v}$ is principal, the former holds if and only if $n=2$ and $b\in\mathcal{O}_{v}^{\times}$, and the latter holds if and only if $n=1$ and $b\notin\mathcal{O}_{v}^{\times}$.
@@
The following facts are used in [4].
@@
Take any valuation ring $\mathcal{O}$ with maximal ideal $\mathcal{M}$. Then
@@
Assume that $a\in\mathcal{O}_{v}\,$, take $\mathcal{O}\in\mathcal{R}$ with maximal ideal $\mathcal{M}$ and denote the valuation associated with $\mathcal{O}$ by $w$.
1) The following holds:
$a\mathcal{M}:(a\mathcal{M})^{n}\>=\>\left\{\begin{array}[]{ll}(a\mathcal{M})^{n-1}&\mbox{if $\mathcal{M}$ is a principal $\mathcal{O}$-ideal, }\\ (a\mathcal{O})^{n-1}=a^{n-1}\mathcal{O}&\mbox{if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal.}\end{array}\right.$
@@
It remains to show that $J\mathcal{O}(IJ)^{\diamondsuit}=J^{\diamondsuit}$ is equal to $(a\mathcal{O})^{n-1}$ if $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. In this case, $J=(a\mathcal{M})^{n-1}=a^{n-1}\mathcal{M}$, and 00 is the infimum of $w\mathcal{M}$, not contained in $w\mathcal{M}$. Therefore, $wa^{n-1}\notin wa^{n-1}+w\mathcal{M}=wa^{n-1}\mathcal{M}$ is the infimum of $wa^{n-1}\mathcal{M}$. Hence by definition, $J^{\diamondsuit}=\{c\mid wc\geq wa^{n-1}\}=a^{n-1}\mathcal{O}=(a\mathcal{O})^{n-1}$. This completes the proof of part 1).
2): Assume first that $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=\mathcal{M}_{v}\,$. Then $a\mathcal{M}/(a\mathcal{M})^{n}$ is nontrivial, whence $a\mathcal{M}\neq(a\mathcal{M})^{n}$, which by part 3) of Lemma 3.20 means that $a\in\mathcal{M}$ or $\mathcal{M}$ is a principal $\mathcal{O}$-ideal. If the latter holds, then by part 1), $a\mathcal{M}:(a\mathcal{M})^{n}=(a\mathcal{M})^{n-1}$. Since $a\mathcal{M}\subseteq\mathcal{M}_{v}\,$, this is equal to $\mathcal{M}_{v}$ if and only if $n-1=1$, $\mathcal{M}=\mathcal{M}_{v}\,$, and $va=0$, where we have used parts 1) and 3) of Lemma 3.20. On the other hand, if the latter does not hold, then $a\in\mathcal{M}$ and by part 1), $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=(a\mathcal{O})^{n-1}\subseteq a\mathcal{O}$. The latter is contained in $\mathcal{M}$ but is not equal to $\mathcal{M}$ since $\mathcal{M}$ is a nonprincipal $\mathcal{O}$-ideal. Thus $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=a\mathcal{M}:(a\mathcal{M})^{n}\subsetneq\mathcal{M}\subseteq\mathcal{M}_{v}$ and consequently, $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}\neq\mathcal{M}_{v}\,$.
For the converse, assume that $n=2$, $a\notin\mathcal{M}_{v}$ and $\mathcal{M}=\mathcal{M}_{v}$ is a principal $\mathcal{O}_{v}$-ideal. Then by part 1) together with part 1) of Lemma 3.20, $\mbox{\rm ann}\,a\mathcal{M}/(a\mathcal{M})^{n}=\mbox{\rm ann}\,a\mathcal{M}_{v}/(a\mathcal{M}_{v})^{n}=(a\mathcal{M}_{v})^{n-1}=a\mathcal{M}_{v}=\mathcal{M}_{v}\,$. ∎